Differentiate f(x) = (x³+4) / (4x- 5 ) using quotient rule.
1
Expert's answer
2020-09-14T11:14:08-0400

f'(x) = ((x³+4)'(4x- 5 )-(x³+4)(4x- 5 )') / (4x- 5 )2 =

(3x2(4x- 5 )-(x³+4)4) / (4x- 5 )2 = (12x3-15x2 )-4x³-16) / (4x- 5 )2 =

(8x3-15x2-16) / (4x- 5 )2 .